Problem 3.12 A borrower borrows $50,000 for 5 years at an annual loan rate of 7.5%; he pays $6,000 at the end for each of the first 2 years and then pays K in the remaining 3 years. Find K.
Answer #1

PV = 50000
Rate = 7.5%

PV = 6000/(1+7.5%) + 6000/(1+7.5%)2 + K/(1+7.5%)3 + K/(1+7.5%)4 + K/(1+7.5%)5​​​​​​​
50000 = 10773.99 + K(1/1.075^3 + 1/1.075^4 + 1/1.075^5)
50000 -10773.99 = K*2.25302
K = 17431.57


Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well